# Towards comprehensive cellular characterisation of H&E slides

### v2

@@ Abstract
Cell detection, segmentation and classification are essential for analyzing tumor microenvironments (TME) on hematoxylin and eosin (H&E) slides. Existing methods suffer from poor performance on understudied cell types (rare or not present in public datasets) and limited cross-domain generalization. To address these shortcomings, we introduce HistoPLUS, a state-of-the-art model for cell analysis, trained on a novel curated pan-cancer dataset of 108,722 nuclei covering 13 cell types. In external validation across 4 independent cohorts, HistoPLUS outperforms current state-of-the-art models in detection quality by 5.2% and overall F1 classification score by 23.7%, while using 5x fewer parameters. Notably, HistoPLUS unlocks the study of 7 understudied cell types and brings significant improvements on 8 of 13 cell types. Moreover, we show that HistoPLUS robustly transfers to 2 oncology indications unseen during training. To support broader TME biomarker research, we release the model weights and inference code at https://github.com/owkin/histoplus/.
@@ 1 Introduction
Cell detection, segmentation and classification (CellDeSC) in histological images form the foundation of modern computational pathology, enabling quantitative analysis of the tumor microenvironment (TME) for applications in patient stratification [12, 1] and personalized medicine [26, 20, 56]. Beyond cell identification, these techniques reveal spatial relationships between tumor cells, immune cells, and stromal compartments that correlate with disease outcomes [7, 5] and treatment responses [45, 51]. The spatial organization and density of specific cell populations – such as tumor-associated neutrophils (linked to immunosuppression [39]) or cancer-associated fibroblasts (implicated in therapy resistance [14, 53, 36]) – have become increasingly important in clinical research and trials [3, 28]. While immunohistochemical methods provide targeted molecular information, H&E staining remains predominant in clinical practice due to its accessibility, cost-effectiveness, and rich morphological detail [19].
@@
Recent pathology foundation models (PFMs) pre-trained via self-supervision [54, 44, 8] promise to mitigate these limitations by leveraging unlabeled histology data to learn transferable representations, reducing reliance on scarce expert annotations. Models like UNI [9], Phikon [16, 17], and H-Optimus-0 [46] have advanced patch classification and molecular prediction [2], but their applicability to CellDeSC – particularly for understudied subtypes, i.e. non-lymphoid immune cells (eosinophils, neutrophils and macrophages), plasmocytes, stromal cells (smooth muscle, endothelial) and rare events (mitotic figures, apoptotic bodies) – has not been thoroughly investigated. Despite evaluations of Hibou-Large [43], UNI, Virchow and GigaPath [48] on cell-level tasks, no benchmark compares PFMs performance rigorously and exhaustively for CellDeSC. This gap is critical, as without systematic benchmarking of PFMs, researchers cannot identify which models maximize performance for CellDeSC tasks – leaving potential performance gains unrealized despite their importance for quantifying the TME.
@@
  • A PFM-integrated model for pan-cancer CellDeSC in H&E, leveraging self-supervised distilled features [17, 46, 18] within a CellViT architecture that achieves state-of-the-art performances on all nuclei identification, including understudied types.
Our model demonstrates superior performance both in cross-validation and on external validation. It paves the way for new analyses of spatial tumor-immune interactions. To facilitate community adoption and further progress in computational pathology, we release our pre-trained CellViT model. This contribution provides the research community with a state-of-the-art CellDeSC model on H&E slides. It aims to accelerate biomarker discovery, ultimately enhancing our ability to translate histological insights into clinically relevant applications.
@@ 2.2 Robust external validation sets are built using a consensus of expert pathologists
Deriving a reliable ground truth for CellDeSC tasks on H&E images is challenging due to substantial inter-annotator variability, even among expert pathologists [30]. Key sources of variation include cell type attribution and annotation completeness. To address these challenges, we randomly assigned 3 out of 9 pathologists to annotate each tile and used a consensus framework to create curated, reliable external validation sets.
@@ 2.3 Pathology foundation models improve cell classification
To evaluate the impact of pathology-specific pretraining on CellDeSC, we integrated PFMs as encoders within the CellViT [27] architecture and compared them with the general-purpose Segment Anything Model [32] (SAM) using 3-fold cross-validation stratified by cell type on HistoTRAIN. To further evaluate the performance of each model to generalize, we performed external validation on HistoVAL. For cell detection and segmentation, we used the detection quality (DQ) and segmentation quality (SQ) metrics. For cell classification we computed F1 scores for each cell type (see Evaluation metrics for detection, segmentation and classification in Methods).
@@ 3 Discussion
We present HistoPLUS, a model that advances the state-of-the-art of cell detection, segmentation and classification. Our results demonstrate that pathology-specific foundation models, and diverse training data together significantly increase performance without requiring overly large models, which is crucial for clinical applications.
@@ 4 Methods
Using the tiles presented in the Data collection paragraph, we chose regions of interest using A) clustering methods to ensure tissue diversity and B) active learning-based strategies to maximize the representation of areas with rare cells or ambiguous cells. As the inference of large CellDeSC is computationally expensive, we used tile-level classifiers performances (simple predictors estimating the presence/absence of a given cell type within a tile) as proxies for estimating the classification performance of CellDeSC models. It allowed us to leverage active learning strategies (iteratively refine classifiers thanks to newly collected annotations) without the compute and time cost of fitting large CellDeSC models at each iteration. We defined 3 different active learning methods to select tiles within candidate datasets; our goal was to not only generate tiles for annotation, but also to evaluate the performance of tile selection strategies and the added value of rare cell enrichment and uncertainty sampling.
@@
Self-supervised learning. Self-supervised learning (SSL) enables models like ViTs to learn robust feature representations from unlabeled data by solving pretext tasks, eliminating the need for costly manual annotations. In computational pathology, SSL frameworks like iBOT [54] (masked image modeling with online tokenization, where patches are reconstructed after random masking), DINO (self-distillation with no labels, using a teacher-student framework to match feature distributions across augmented views), and DINOv2 [44] (a scalable extension of DINO with improved regularization, larger datasets, and efficient data pipelines) have become pivotal for pretrained foundation models. While iBOT focuses on local feature consistency through masked prediction, DINO and DINOv2 emphasize global representation alignment, with DINOv2 further optimizing stability and scalability for large-scale training. These methods leverage vision transformer architectures to capture hierarchical and semantically rich features from histopathology patches, which can later be fine-tuned for downstream tasks like CellDeSC. UNI and UNI2 were trained using DINOv2, enabling strong generalization across diverse tissue types. Virchow and Virchow2G were trained using DINOv2. Phikon and PhikonV2 were trained respectively using iBOT and DINOv2.
@@ 6 Code Availability
An implementation of the NuClick model is available at https://github.com/mostafajahanifar/nuclick_torch/. The packaged models and the code used to infer the segmentation models can be found at https://github.com/owkin/histoplus, under a Creative Commons Attribution 4.0 International (CC-BY-4.0) License. The repositories referenced above contain license files and details of usage permissions. Any reuse of third-party software, including the NuClick implementation and the models used, complies with their respective licenses.

### v3

@@ Abstract
Cell detection, segmentation and classification are essential for analyzing tumor microenvironments (TME) on hematoxylin and eosin (H&E) slides. Existing methods suffer from poor performance on understudied cell types (rare or not present in public datasets) and limited cross-domain generalization. To address these shortcomings, we introduce HistoPLUS, a state-of-the-art model for cell analysis, trained on a novel curated pan-cancer dataset of 108,722 nuclei covering 13 cell types. In external validation across 4 independent cohorts, HistoPLUS outperforms current state-of-the-art models in detection quality by 5.2% and overall F1 classification score by 23.7%, while using 5x fewer parameters. Notably, HistoPLUS unlocks the study of 7 understudied cell types and brings significant improvements on 8 of 13 cell types. Moreover, we show that HistoPLUS robustly transfers to 2 oncology indications unseen during training. To support broader TME biomarker research, we release the model weights and inference code (see Code Availability).
@@ 1 Introduction
Cell detection, segmentation and classification in histological images form the foundation of modern computational pathology, enabling quantitative analysis of the tumor microenvironment (TME) for applications in patient stratification [12, 1] and personalized medicine [26, 20, 56]. Beyond cell identification, these techniques reveal spatial relationships between tumor cells, immune cells, and stromal compartments that correlate with disease outcomes [7, 5] and treatment responses [45, 51]. The spatial organization and density of specific cell populations – such as tumor-associated neutrophils (linked to immunosuppression [39]) or cancer-associated fibroblasts (implicated in therapy resistance [14, 53, 36]) – have become increasingly important in clinical research and trials [3, 28]. While immunohistochemical methods provide targeted molecular information, H&E staining remains predominant in clinical practice due to its accessibility, cost-effectiveness, and rich morphological detail [19].
@@
Recent pathology foundation models (PFMs) pre-trained via self-supervision [54, 44, 8] promise to mitigate these limitations by leveraging unlabeled histology data to learn transferable representations, reducing reliance on scarce expert annotations. Models like UNI [9], Phikon [16, 17], and H-Optimus-0 [46] have advanced patch classification and molecular prediction [2], but their applicability to cell detection, segmentation and classification – particularly for understudied subtypes, i.e. non-lymphoid immune cells (eosinophils, neutrophils and macrophages), plasmocytes, stromal cells (smooth muscle, endothelial) and rare events (mitotic figures, apoptotic bodies) – has not been thoroughly investigated. Despite evaluations of Hibou-Large [43], UNI, Virchow and GigaPath [48] on cell-level tasks, no benchmark compares PFMs performance rigorously and exhaustively for simultaneous cell detection, segmentation and classification tasks. This gap is critical, as without systematic benchmarking of PFMs, researchers cannot identify which models maximize performance for these tasks – leaving potential performance gains unrealized despite their importance for quantifying the TME.
@@
  • A PFM-integrated model for pan-cancer cell detection, segmentation and classification in H&E, leveraging self-supervised distilled features [17, 46, 18] within a CellViT architecture that achieves state-of-the-art performances on all nuclei identification, including understudied types.
Our model demonstrates superior performance both in cross-validation and on external validation. It paves the way for new analyses of spatial tumor-immune interactions. To facilitate community adoption and further progress in computational pathology, we release our pre-trained CellViT model. This contribution provides the research community with a state-of-the-art cell detection, segmentation and classification model on H&E slides. It aims to accelerate biomarker discovery, ultimately enhancing our ability to translate histological insights into clinically relevant applications.
@@ 2.2 Robust external validation sets are built using a consensus of expert pathologists
Deriving a reliable ground truth for cell detection, segmentation and classification tasks on H&E images is challenging due to substantial inter-annotator variability, even among expert pathologists [30]. Key sources of variation include cell type attribution and annotation completeness. To address these challenges, we randomly assigned 3 out of 9 pathologists to annotate each tile and used a consensus framework to create curated, reliable external validation sets.
@@ 2.3 Pathology foundation models improve cell classification
To evaluate the impact of pathology-specific pretraining on cell detection, segmentation and classification, we integrated PFMs as encoders within the CellViT [27] architecture and compared them with the general-purpose Segment Anything Model [32] (SAM) using 3-fold cross-validation stratified by cell type on HistoTRAIN. To further evaluate the performance of each model to generalize, we performed external validation on HistoVAL. For cell detection and segmentation, we used the detection quality (DQ) and segmentation quality (SQ) metrics. For cell classification we computed F1 scores for each cell type (see Evaluation metrics for detection, segmentation and classification in Methods).
@@ 3 Discussion
We present HistoPLUS, a model that advances the state-of-the-art of simultaneous cell detection, segmentation and classification. Our results demonstrate that pathology-specific foundation models, and diverse training data together significantly increase performance without requiring overly large models, which is crucial for clinical applications.
@@ 4 Methods
Using the tiles presented in the Data collection paragraph, we chose regions of interest using A) clustering methods to ensure tissue diversity and B) active learning-based strategies to maximize the representation of areas with rare cells or ambiguous cells. As the inference of large cell detection, segmentation and classification is computationally expensive, we used tile-level classifiers performances (simple predictors estimating the presence/absence of a given cell type within a tile) as proxies for estimating the classification performance of cell detection, segmentation and classification models. It allowed us to leverage active learning strategies (iteratively refine classifiers thanks to newly collected annotations) without the compute and time cost of fitting large cell detection, segmentation and classification models at each iteration. We defined 3 different active learning methods to select tiles within candidate datasets; our goal was to not only generate tiles for annotation, but also to evaluate the performance of tile selection strategies and the added value of rare cell enrichment and uncertainty sampling.
@@
Self-supervised learning. Self-supervised learning (SSL) enables models like ViTs to learn robust feature representations from unlabeled data by solving pretext tasks, eliminating the need for costly manual annotations. In computational pathology, SSL frameworks like iBOT [54] (masked image modeling with online tokenization, where patches are reconstructed after random masking), DINO (self-distillation with no labels, using a teacher-student framework to match feature distributions across augmented views), and DINOv2 [44] (a scalable extension of DINO with improved regularization, larger datasets, and efficient data pipelines) have become pivotal for pretrained foundation models. While iBOT focuses on local feature consistency through masked prediction, DINO and DINOv2 emphasize global representation alignment, with DINOv2 further optimizing stability and scalability for large-scale training. These methods leverage vision transformer architectures to capture hierarchical and semantically rich features from histopathology patches, which can later be fine-tuned for downstream tasks like cell detection, segmentation and classification. UNI and UNI2 were trained using DINOv2, enabling strong generalization across diverse tissue types. Virchow and Virchow2G were trained using DINOv2. Phikon and PhikonV2 were trained respectively using iBOT and DINOv2.
@@ 6 Code Availability
To support broader TME biomarker research, we release the model weights and inference code at https://github.com/owkin/histoplus/, under a Creative Commons Attribution 4.0 International (CC-BY-4.0) License. An implementation of the NuClick model is available at https://github.com/mostafajahanifar/nuclick_torch/. The repositories referenced above contain license files and details of usage permissions. Any reuse of third-party software, including the NuClick implementation and the models used, complies with their respective licenses.